A principal of $7500 is invested in an account paying an annual rate of 7%. Find the amount in the account after 3 years if the account is compounded semiannually, quarterly and monthly.

Answer by checkley79(3341)

P(1+I)^T
7500(1+.07/2)^3*2
7500(1.035)^6
7500*1.229=$9,219.41 ANS FOR SEMIANNUAL INTEREST.
7500(1+.07/4)^3*4
7500(1.0175)^12
7500*1.2314=$9,235.79 ANS FOR QUARTERLY INTEREST.
7500(1+.07/12)^3*12
7500(1+.005833)^36
7500*1.2329=$9,246.83 ANS FOR MONTHLY INTEREST.